The electric potential at x=3.00 m is 120 V, and the electric potential at x=5.00 m is 190 V. What is the x-component of the electric in this region, assuming the field is uniform?
1
Expert's answer
2020-10-19T13:20:00-0400

The x-component of the electric field= -change in electric potential "\div" change in position"=-\frac{190V-120V}{5m-3m}=-\frac{70V}{2m}=-35\frac{V}{m}=-35\frac{N}{C}"